### v1

Abstract
The fundamental information-theoretic limits of covert, or low probability of detection/intercept (LPD/LPI), communication have been extensively studied for over a decade, resulting in the square root law (SRL): only $L\sqrt{n}$ covert bits can be reliably transmitted over time-bandwidth product $n$, for constant $L>0$. Transmitting more either results in detection or decoding errors. The SRL imposes significant constraints on hardware realization of mathematically-guaranteed covert communication. Indeed, they preclude using standard link maintenance operations that are taken for granted in non-covert communication. Thus, experimental validation of covert communication is underexplored: to date, only two experimental studies of SRL-based covert communication are available, both focusing on optical channels. Here, we report a demonstration of provably-secure covert radio-frequency (RF) communication using software-defined radios (SDRs). This validates theoretical predictions, opens practical avenues for implementing covert communication systems, and raises further research questions.
I Introduction
Covert, or low probability of detection/intercept (LPD/LPI) communication, allows transmitting messages without alerting an adversary [1, 2, 3, 4]. This mode of information security is highly desirable and contrasts the traditional cryptographic [5] and information-theoretic secrecy [6] methods that prevent adversaries’ access to the transmission’s content, but not its detection. Careful waveform design and spread spectrum techniques are often employed in practice to reduce adversaries’ signal-to-noise ratio (SNR) below the noise floor [7, Pt. 1, Ch. 5]. However, guaranteeing covertness requires intricate mathematical analysis which yields the square root law (SRL): only $B(n)=L\sqrt{n}$ covert bits can be reliably transmitted over $n$ channel uses [1, 2, 3, 4]. The channel-dependent constant $L>0$ is called covert capacity and $n=TW$ is typically the time-bandwidth product for the transmission. Notably, the Shannon capacity [8] of this channel is zero, since $\lim_{n\to\infty}\frac{B(n)}{n}=0$, This is because the adversary’s task in covert communication involves learning just one bit of information (whether transmitter is on or not) versus $\mathcal{O}(n)$ bits of transmitted data in traditional secure communication. Nevertheless, a significant number of such mathematically-guaranteed covert bits can still be transmitted in a finite amount of time.
The discovery of the SRL in [1, 2] resulted in an explosion of research by the communication and information theory communities overviewed in a tutorial [3] and a detailed survey [4]. This included characterization of capacity $L$ for additive white Gaussian noise (AWGN) and discrete memoryless channels (DMCs) [9, 10, 11, 12], covert networks [13, 14, 15, 16, 17, 18], quantum aspects of covert communication [19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30], and many other directions. Notwithstanding this progress on fundamental theory, experimental SRL-based covert communication remains underexplored, with only two published works, both focused on optical channels [20, 31]. This paper addresses this gap by providing the first, to our knowledge, experimental validation of the SRL on radio frequency (RF) channels.
We implement and evaluate a covert communication protocol using software-defined radio (SDR). Our implementation uses USRP X310 SDR units deployed on the ORBIT testbed [32, 33], as detailed in Section III-A, enabling controlled and reproducible experiments. We employ quadrature phase-shift keying (QPSK) with a Gaussian pulse shaping filter. This controls temporal and spectral symbol leakage while mitigating timing jitter. We implement an experimental framework based on a synthetic Gaussian noise source within a wired, shielded network that provides environmental control and supports reproducibility.
The strict output power constraint of SRL-based covert communication precludes using standard handshake, carrier synchronization, channel state estimation, and other link maintenance operations that are taken for granted in non-covert communication. This necessitates using precise timing sources, such as atomic clocks. Furthermore, the finite resolution of SDRs’ digital-to-analog and analog-to-digital converters (DACs and ADCs) imposes a lower limit on the amplitude of the transmitted symbol. Thus, to meet the SRL, we must use a “sparse coding” strategy, selecting a random subset of available channel uses that is secretly shared with the receiver in advance. This sparsity, in turn, renders the standard phase synchronization approaches that rely on periodic pilot signal transmission unusable. Therefore, we include a pilot signal with every transmitted symbol, which is done rarely if at all in non-covert communications. This, coupled with precise time synchronization, enables our SDR implementation to transmit a significant number of bits with a covertness guarantee.
The SRL governs, arguably, the worst case scenario. As detailed in Section II, it assumes that the characteristics of the transmitter-adversary channel (noise power variance and transmission loss), the exact time and frequency band of potential transmission, and details of the transceiver system design are known to the adversary. The adversary’s only limitations are: 1) inability to control the random channel noise; and, 2) lack of access to the secret shared between transmitter and receiver prior to the transmission. This is the model implemented in our system. While conservative, it provides a high level of security to unforeseen adversarial technological surprises. Indeed, relaxing these assumptions can result in significant gains in system performance. For example, an adversary’s uncertainty of transmission time/frequency yields a multiplicative improvement to covert capacity $L$ [34, 35, 36], while uncertainty in noise power level can lead to a linear law: $\mathcal{O}(n)$ covert bits reliably transmissible over $n$ channel uses [37, 38]. We defer validation of these to future work. Additionally, our experiment motivates a systematic study of the impact of adversary hardware limitations such as sampling timing jitter and receiver bandwidth constraints. Furthermore, employing more efficient modulation and coding schemes (such as [39]), as well as minimizing the amount of pre-shared secret (see [9]) are avenues for future exploration.
This paper is organized as follows: Section II develops the theoretical foundation for radio-frequency covert communication, beginning with the discrete-time AWGN channel model and deriving a sparse-coded QPSK transmission scheme that satisfies the SRL. Section III details our SDR implementation on the ORBIT testbed, describes the experimental procedure, and presents measurement results that validate the theory. Section IV interprets these findings, highlighting practical design challenges and their implications for real-world systems, and outlines open problems and promising directions for extending this work. Finally, Appendices A-C provide supporting derivations for the covertness criterion, phase-offset estimation, and SNR estimation used in the experiments.
II Analysis of Practical Covert Communications
II-A Channel Model
We employ a discrete-time AWGN block channel model in Fig. 1. Alice uses a discrete-time $n_{s}$-length complex-valued pulse shape $\vec{u}(x)\in\mathbb{C}^{n_{s}}$ to modulate data in $x\in\mathcal{X}$, where the modulation constellation $\mathcal{X}$ is defined in Section II-B. This pulse is transmitted over $n_{s}$ channel uses to legitimate receiver Bob and adversary (warden) Willie. Bob receives $h_{b}\vec{u}(x)+\vec{z}^{(b)}$ while Willie observes $h_{w}\vec{u}(x)+\vec{z}^{(w)}$. The respective complex-valued channel gains $h_{b}=a_{b}e^{j\theta_{b}}$ and $h_{w}=a_{w}e^{j\theta_{w}}$ contain attenuation $a_{b},a_{w}>0$ and phase $\theta_{b},\theta_{w}\in[0,2\pi]$ components. We assume channel coherence time longer than the pulse duration, thus, channel gains $h_{b}$ and $h_{w}$ remain constant at least over the $n_{s}$ AWGN channel uses. For each pulse, attenuation is arbitrary, while phase is assumed to be selected uniformly at random from $0,2\pi]$. Each symbol in the transmitted pulse $\vec{u}(x)$ is corrupted independently by AWGN: $\vec{z}^{(b)}$ and $\vec{z}^{(w)}$ are instances of independent and identically distributed (i.i.d.) circularly-symmetric complex Gaussian random vectors $\vec{Z}^{(b)}\sim\mathcal{CN}\left(\vec{0},2\sigma_{b}^{2}\mathbf{I}_{n_{s}}\right)$ and $\vec{Z}^{(w)}\sim\mathcal{CN}\left(\vec{0},2\sigma_{w}^{2}\mathbf{I}_{n_{s}}\right)$, where $\mathbf{I}_{n_{s}}$ is an $n_{s}\times n_{s}$ identity matrix.
The pulse-to-pulse variation in channel gain represents the dynamic operating environment due to the mobility of Alice, Bob, and Willie, as well as the surrounding reflecting objects. Assigning a suitable probability distribution (e.g., Rayleigh, Rice, or Nakagami) to attenuation coefficients $a_{b}$ and $a_{w}$ (along with uniform phase distribution) yields a corresponding block fading model (see [40, Ch. 2]). Here phases $\theta_{b}$ and $\theta_{w}$ are random for each pulse slot, which captures not only environment dynamics but also the drift of independent oscillators. However, we assume for simplicity that attenuation is constant on all channels. This models a scenario where, say, worst-case path-loss figures are employed. We defer the analysis of fully-dynamic channels to future work.
Figure 1: Discrete-time AWGN block broadcast channel. Alice modulates $x\in\mathcal{X}$ into a complex-valued pulse shape $\vec{u}(x)\in\mathbb{C}^{n_{s}}$ and transmits over $n_{s}$ uses of a channel corrupted by independent AWGN at legitimate receiver Bob and adversary (warden) Willie. Bob receives $h_{b}\vec{u}(x)+\vec{z}^{(b)}$ while Willie observes $h_{w}\vec{u}(x)+\vec{z}^{(w)}$, where the channel gain $h_{r}=a_{r}e^{j\theta_{r}}$ from Alice to the receiver $r\in\{b,w\}$ contains attenuation $a_{r}>0$ and a random phase $\theta_{r}$. We model noise $\vec{z}^{(r)}$ by circularly-symmetric complex Gaussian random vectors.
II-B Practical Covert System Design
Figure 2: System model for covert communication. Alice either encodes, modulates, and transmits a message in $n_{p}$ pulses over $n=n_{p}n_{s}$ uses of the AWGN channel, or she remains silent. Here, $\vec{x}\in\left(\mathcal{X}\cup\{0\}\right)^{n_{p}}$ is the vector of QPSK and zero symbols, with the relative location of each selected using $\vec{t}$. After modulation, $\vec{u}(\vec{x})\in\mathbb{C}^{n}$ is the input pulse-shaped encoded signal from Alice. AWGN channel is characterized by additive noise $\vec{z}^{(w)}\in\mathbb{C}^{n}$ and channel gain $\vec{h}_{b},\vec{h}_{w}\in\mathbb{C}^{n_{p}}$, as depicted in Fig. 1. Here, each element of $\vec{h}_{b},\vec{h}_{w}$ applies uniformly to one $n_{s}$ sample pulse slot. Then, $\vec{y}(\vec{x})\in\mathbb{C}^{n}$ (resp. $\vec{w}(\vec{x})\in\mathbb{C}^{n}$) is the output at Bob (resp. Willie). Willie’s goal is to determine whether Alice is transmitting. Bob and Alice use a pre-shared secret $(\vec{s},\vec{t})$ to prevent this while communicating reliably.
The square root law (SRL) governs covert communication over the AWGN channel in Fig. 1: one can transmit $B(n)=L\sqrt{n}$ covert bits reliably in $n$ uses of such channel [1, 2], where $L>0$ is the covert capacity (derived in [9, 10]) and $n=TW$ is the transmission time-bandwidth product. These results and much of the follow-up employ a simplified scenario with $n_{s}=1$, $a_{b}=a_{w}=1$, and $\theta_{b}=\theta_{w}=0$. However, practical radio systems require pulse-shaping to efficiently use available bandwidth and to mitigate inter-symbol interference and timing jitter. Furthermore, analysis of the fundamental limits often assumes that Alice can generate inputs at arbitrarily low power. In practice, DACs and ADCs have finite resolution, which results in a lower limit on output signal power. We address this next, and then discuss modulation and pulse shape. Our overall system model is depicted in Fig. 2.
1 Sparse Coding
Divide $n$ available channel uses into $n_{p}$ pulses, with shape given by the $n_{s}$-length pulse-shape vector $\vec{u}(x)$ discussed next, $x\in\mathcal{X}$. Then, $n=n_{p}n_{s}$, with $n_{p}\in\mathcal{O}(n)$. The minimum output power limit implies the norm $\|\vec{u}(x)\|=c$ for all $x\in\mathcal{X}$. We use sparse coding to ensure covertness: before transmission, Alice and Bob randomly generate and secretly share an $n_{p}$-length sequence $\vec{t}$ of i.i.d. samples from the Bernoulli distribution $p(t_{i})=\{1-\alpha_{n}\text{ if }t_{i}=0;\alpha_{n}\text{ if }t_{i}=1\}$, where the probability of transmitting in a given slot $\alpha_{n}\in\mathcal{O}\left(1/\sqrt{n}\right)$ follows the SRL [1, 2]. The number $n_{t}=\sum_{i=1}^{n_{p}}t_{i}$ of selected pulse slots is a random variable with mean $\alpha_{n}n_{p}$. It is also the length of the transmitted message $\vec{x}_{n_{t}}$, in symbols. Alice is quiet in the slots that are not selected, that is, when $t_{i}=0$, she inputs $\vec{0}$. Denoting by $x=0$ this “innocent” symbol, the transmitted symbol vector $\vec{x}\in\left(\mathcal{X}\cup\{0\}\right)^{n_{p}}$ contains $n_{t}$ symbols from $\mathcal{X}$ and $n_{p}-n_{t}$ innocent symbols. This yields Alice’s sparse-coded covert input $\vec{u}(\vec{x})\in\mathbb{C}^{n}$, containing mostly innocent symbols. We derive $\alpha_{n}$ in Section II-D. We assume that Bob and Alice have a reference clock that allows them to synchronize the start of transmission.
2 Modulation and Pulse Shape
The channel model in Section II-A includes random pulse-to-pulse phase variation. On-off keying (OOK) modulation is resilient to this, as it uses energy pulses. However, OOK is power-inefficient; matching the per-symbol bit error rate (BER) of binary and quadrature phase-shift keying (BPSK and QPSK) schemes requires transmission of twice the power, which, in turn, doubles Willie’s received SNR and harms covertness. We discuss the relationship between SNR and covert signal construction further in Section II-D. We choose QPSK modulation since it doubles the number of degrees of freedom relative to BPSK, further increasing the achievable covert communication rate.
Phase variations cause errors in phase-keyed systems, which are typically mitigated by periodic pilot signals. However, the SRL does not allow periodic transmissions. Furthermore, the sparse coding scheme results in time between pulses scaling as $\mathcal{O}(\sqrt{n})$ on average, rendering phases uncorrelated for large enough $n$ when channel coherence time is reasonably assumed to be finite. Thus, we include a fixed $n_{s}^{(p)}$-length pilot segment $\vec{c}_{p}$ in every transmitted pulse. For long enough channel coherence time, the pilot segment allows estimation of the phase to correct the $n_{s}^{(q)}$-length QPSK-modulated data segment $\vec{c}_{q}(x)=e^{\frac{j\pi(2x-1)}{4}}\vec{c}_{q}$, $x\in\mathcal{X}$, where $\mathcal{X}=\{1,\ldots,4\}$ and each $x\in\mathcal{X}$ represented by two bits. Further, since the relative phase between pilot and QPSK pulse is what needs to be estimated, Alice randomly selects $\theta$ from a uniform distribution over $0,2\pi)$ and applies to both pilot and QPSK pulse in each transmitted pulse. This improves covertness by ensuring that the relative phase between transmissions is randomized without any additional cost to decoding reliability, since Bob needs to estimate the phase anyway. Hence, the transmitted pulse is given by $\vec{u}(x)=\left(e^{j\theta}\vec{c}_{p},e^{j\theta}\vec{c}_{q}(x)\right)$ as depicted in Fig. 3, with length $n_{s}=n_{s}^{(p)}+n_{s}^{(q)}$ and norm $\|\vec{u}(x)\|\triangleq c$ for all $x\in\mathcal{X}$. Since the pilot requires increased power output per pulse over standard QPSK, this scheme increases Willie’s received SNR based the chosen magnitude of the pilot. Thus, there is a trade-off between phase estimation and covertness of our signal that we optimize in Section III-C.
Figure 3: Pulse shape $\vec{u}(1)$. A strictly positive pilot segment $\vec{c_{p}}$ precedes the data segment $\vec{c_{q}}$, whose in-phase and quadrature components form the QPSK symbol $e^{\frac{j\pi}{4}}$ encoding $x=1$. The projections of the complex waveform onto the real and imaginary axes are overlaid with dotted red and blue lines, respectively, and the discrete time instants at which the Gaussian envelope is sampled for transmission are marked by $\star$.
Sparse coding has negligible inter-symbol interference. Instead covertness demands energy confinement in time and frequency domains. Thus, we shape both $\vec{c}_{p}$ and $\vec{c}_{q}$ by a discrete-time Gaussian envelope:
$\vec{g}_{k}[m]=\exp\left({-\frac{\left(m-n_{s}^{(k)}/2\right)^{2}}{2\sigma_{k}^{2}}}\right),\quad\left.\begin{aligned} m&=0,\dots,n_{s}^{(k)}-1,\\ k&\in\{p,q\}\end{aligned}\right$. (1)
and define
$\vec{c}_{k}\triangleq c_{k}\frac{\vec{g}_{k}}{\|\vec{g}_{k}\|},\quad\left.\begin{aligned} k&\in\{p,q\}\end{aligned}\right$. (2)
where $\sigma_{p}$, $\sigma_{q}$, $c_{p}$, and $c_{q}$ are the corresponding pulse width and magnitude parameters. We note that more complicated pulse shapes with lower bandwidth under the same power and time constraints exist, e.g., discrete prolate spheroidal (Slepian) sequences [41, Sec. 3.1] but we defer using them to future work. In Section III-C, we optimize $n_{s}^{(p)}$, $n_{s}^{(q)}$, $c_{p}$, and $c_{q}$. Thus, we choose the largest $\sigma_{k}$ satisfying
$\frac{\sum_{m=0}^{n_{s}^{(k)}-1}\vec{c}_{k}[m]}{\sum_{m=-\infty}^{\infty}\vec{c}_{k}[m]}\;\geq\;0.999,\qquad k\in\{p,q\},$ (3)
which ensures that at least $99.9\%$ of each pulse’s energy is confined to its designated slot.
II-C Reliability Analysis
Bob has $\vec{t}$ and, thus, can discard the slots that are not used for transmission (that is, those slots corresponding to $x=0$ in $\vec{x}$ when Alice “transmits” $\vec{0}$). Thus, Bob only examines $n_{t}$ pulses in $\vec{y}(\vec{x}_{n_{t}})$. When Alice transmits a symbol $x\in\mathcal{X}$ from $\vec{x}_{n_{t}}$ in each of the $n_{t}$ selected pulse slots, Bob receives the $n_{s}$-sample vector $\vec{y}=\left(\vec{y}_{p},\vec{y}_{q}\right)$ containing two segments:
$\vec{y}_{p}$ $=a_{b}e^{j\theta_{b}}\vec{c}_{p}+\vec{z}^{(b)}_{p}$ (4)
$\vec{y}_{q}(x)$ $=a_{b}e^{\frac{j\pi(2x-1)}{4}+\theta_{b}}\vec{c}_{q}+\vec{z}^{(b)}_{q}$ (5)
where $a_{b}$ is a constant loss, $\theta_{b}$ is unknown phase offset uniformly-distributed from $0,2\pi]$, and $\vec{z}^{(b)}=(\vec{z}^{(b)}_{p},\vec{z}^{(b)}_{q})$ is complex-valued circularly-symmetric AWGN, per Section II-A. Bob uses $\vec{y}_{p}$ to obtain the estimate $\hat{\theta}_{b}$ of $\theta_{b}$ (see Appendix B), and correct the phase of the data segment by applying $\vec{y}_{q}(x)e^{-j\hat{\theta}_{b}}$. Bob decodes the QPSK symbol by first projecting the phase-aligned data segment onto the pulse shape,
$r_{\mathrm{I}}$ $=\left\langle\Re\{\vec{y}_{q}(x)\,e^{-j\hat{\theta}_{b}}\},\vec{c}_{q}\right\rangle,$ $r_{\mathrm{Q}}$ $=\left\langle\Im\{\vec{y}_{q}(x)\,e^{-j\hat{\theta}_{b}}\},\vec{c}_{q}\right\rangle.$ (6)
This yields two BPSK symbols. Hard decisions $\hat{\tilde{b}}_{1}=\operatorname{sgn}(r_{\mathrm{I}})$ and $\hat{\tilde{b}}_{0}=\operatorname{sgn}(r_{\mathrm{Q}})$, followed by $\hat{b}_{k}=(1-\hat{\tilde{b}}_{k})/2$, $k\in\{0,1\}$, extract the two bits. Since AWGN is symmetric, this hard-decision scheme induces two binary symmetric channels (BSCs) with probability of error:
$p_{e,\mathrm{bsc}}^{(b)}\triangleq\Pr(\hat{b}_{k}=0\mid b_{k}=1)=\Pr(\hat{b}_{k}=1\mid b_{k}=0).$ (7)
For sufficiently large $n_{p}$, Alice and Bob can use an error correction code (ECC) [42] on the approximately $\alpha_{n}n_{p}$-long subset of pulse slots $\{v:t_{v}=1\}$. This allows reliable transmission of
$B_{\rm bsc}(n)$ $=2n_{t}C_{\rm bsc}\approx 2\alpha_{n}n_{p}C_{\rm bsc}$ (8)
bits in $n$ channel uses on average, with $C_{\rm bsc}\triangleq 1-h_{2}\left(p_{e,\rm{bsc}}^{(b)}\right)$, where $h_{2}(p)\triangleq-p\log_{2}(p)-(1-p)\log_{2}(1-p)$ is the binary entropy function and the approximation is due to $n_{t}$ being a random variable.
The ECC structure, which can aid Willie in detection, is eliminated by applying a $2n_{t}$-bit one-time-pad $\vec{s}$ to the encoder output written in binary. Each bit in $\vec{s}$ is selected by Alice and Bob equiprobably at random, resulting in output distribution $\Pr(x=-1)=\Pr(x=1)=\frac{1}{2}$. Pre-shared secret includes $\vec{t}$ and $\vec{s}$.
We require pre-sharing of $\mathcal{O}(\sqrt{n}\log n)$ secret key bits, since $\vec{t}$ and $\vec{s}$ take $n_{t}\log_{2}n_{p}$ and $2n_{t}$ bits to represent, respectively. It is possible to reduce the required number of secret bits to $\mathcal{O}(\sqrt{n})$ at a significant complexity expense [9]. As computation uses more energy than storage, in the context of future portable covert communication systems, the logarithmic cost is worth the simplicity of our scheme.
II-D Hypothesis Testing and Covertness
Willie observes sequence $\vec{w}\in\mathbb{C}^{n}$ and has to decide if Alice is transmitting. Unlike Bob, he does not have access to $\vec{t}$ and $\vec{s}$, but he knows the exact transmission start time, and pulse boundaries. Additionally, we assume that he knows the channel conditions, and other details of transceiver design (including $\vec{c}$ and $\alpha$). He performs a binary hypothesis test between hypotheses $H_{0}$ (no transmission) and $H_{1}$ (transmission). Let distributions $P_{0}^{n}$ and $P_{1}^{n}$ and associated density functions $p_{0}^{n}(\vec{w})$ and $p_{1}^{n}(\vec{w})$ describe the statistics of Willie’s output $\vec{w}$ when Alice is silent ($H_{0}$) and transmitting ($H_{1}$). Assuming non-informative priors [^1] Accounting for arbitrary priors is discussed in [38]. $\Pr(H_{0})=\Pr(H_{1})=\frac{1}{2}$, Willie’s probability of error for an optimal detection scheme is [43, Th. 13.1.1]:
$p_{e}^{(w)}=\frac{1}{2}-\frac{1}{2}\mathcal{V}_{T}(P_{0}^{n},P_{1}^{n}),$ (9)
where $\mathcal{V}_{T}(P_{0}^{n},P_{1}^{n})\triangleq\frac{1}{2}\int_{\mathbb{V}^{n}}\mathop{}\!\mathrm{d}^{2}\vec{w}|p_{0}^{n}(\vec{w})-p_{1}^{n}(\vec{w})|$ is the total variation distance between $P_{0}^{n}$ and $P_{1}^{n}$. We say that the transmission is $\delta$-covert if $p_{e}^{(w)}\geq\frac{1}{2}-\delta$. Total variation distance is mathematically unwieldy, hence, in covert communication literature [1, 2, 9, 10] we often employ Pinsker’s inequality [8, Lemma 11.6.1] to lower bound
$p_{e}^{(w)}$ $\geq\frac{1}{2}-\frac{1}{2\sqrt{2}}\sqrt{D(P_{0}^{n}\|P_{1}^{n})},$ (10)
where $D(P_{0}^{n}\|P_{1}^{n})\triangleq\int_{\mathbb{C}^{n}}\mathop{}\!\mathrm{d}^{2}\vec{w}p_{0}^{n}(\vec{w})\log_{2}\frac{p_{0}^{n}(\vec{w})}{p_{1}^{n}(\vec{w})}$ is the relative entropy of $P_{0}^{n}$ and $P_{1}^{n}$. Thus, instead of (9), it is standard to use (10), and assert that any scheme is $\delta$-covert if $D(P_{0}^{n},P_{1}^{n})\leq\delta_{\rm RE}=8\delta^{2}$. Ultimately, satisfying this relative entropy constraint for a given $\delta$ requires limiting our average output power. While relative entropy between our distributions can be readily bounded, Pinsker’s inequality is loose for many distribution pairs. Thus, we seek a tighter bound than (10), which will allow Alice more average power throughput while still satisfying (9). Inspired by recent work [27] that characterized covert communication over a quantum channel model with fidelity constraint, we employ the classical analog of the fidelity constraint to instead bound (9) via the Cauchy-Schwarz inequality as follows:
$p_{e}^{(w)}$ $\geq\frac{1}{2}-\frac{1}{\sqrt{2}}H(P_{0}^{n},P_{1}^{n})$ (11)
where $H(P_{0}^{n},P_{1}^{n})\triangleq\sqrt{\frac{1}{2}\int_{\mathbb{C}^{n}}\mathop{}\!\mathrm{d}^{2}\vec{w}\left(\sqrt{p_{0}^{n}(\vec{w})}-\sqrt{p_{0}^{n}(\vec{w})}\right)^{2}}$ is the Hellinger distance between $P_{0}^{n}$ and $P_{1}^{n}$.
When not silent, Alice inputs QPSK symbol $x\in\mathcal{X}$ in each pulse slot according to Section II-B. Then, Willie receives $\vec{w}_{p}(x)=h_{w}\vec{u}(x)+\vec{z}^{(w)}$ for $\vec{u}(x)=\left(e^{j\theta}\vec{c}_{p},e^{j\theta}\vec{c}_{q}(x)\right)$, per the model in Fig. 1. Let $r_{p/q}=\frac{\|\vec{c}_{p}\|}{\|\vec{c}_{q}\|}$ be the ratio of magnitudes of the pilot and QPSK pulses, and $c_{q}=\|\vec{c}_{q}\|$ be the magnitude of the QPSK pulse. Appendix A shows that we may ensure $\delta$-covertness of the transmission scheme by setting
$\alpha_{n}$ $=\frac{4\sigma_{w}^{2}}{a_{w}^{2}c_{q}^{2}}\sqrt{\frac{2\log\left(\frac{1}{1-2\delta^{2}}\right)}{n_{p}(1+r_{p/q}^{4})}}$ (12)
$=\left(\mathrm{SNR}\sqrt{1+r_{p/q}^{4}}\right)^{-1}\sqrt{\frac{32\log\left(\frac{1}{1-2\delta^{2}}\right)}{n_{p}}},$ (13)
where $\mathrm{SNR}\triangleq\frac{h_{w}^{2}c_{q}^{2}}{\sigma^{2}_{w}}$ is Willie’s received SNR and the penalty term $\sqrt{1+r_{p/q}^{4}}$ is due to inclusion of the pilot signal for channel phase estimate. Combining (12) with $n_{p}=n/n_{s}$ and (8) yields the SRL scaling of covert throughput.
III Experiment Implementation and Results
III-A System Configuration
We perform our experiments on ORBIT, an open-access radio grid testbed [32]. Our system is depicted in Fig. 4. We connect four Ettus universal software radio peripheral (USRP) X310 SDRs fitted with UBX 160 – corresponding to transmitter Alice, receiver Bob, adversary Willie, and an AWGN generator – into a shared-medium RF star network topology via coaxial cables. This isolates our experiment from others on ORBIT, enables control over the environment, and allows repeatability. We defer a wireless experiment with an organic noise source to future work. The radios are positioned so that Bob and Willie experience approximately equal attenuation from both the AWGN generator and Alice, corresponding to an adverse scenario of Willie being close to Bob. We utilize a single 12.5 MHz channel centered at $f_{c}=915$ MHz.
Each radio has a dedicated enhanced small form-factor pluggable (SFP+) cable connecting it to a high-bandwidth router, and the router to a dedicated control node on the ORBIT grid [33]. SFP+ ports operate in 10 Gbps mode to support the maximum transmission unit (MTU) size of 8 kB. This prevents packet drops from the internal radio buffer overflow during an experiment. We note that even a single packet drop causes catastrophic misalignment of the covert symbols within the transmission, rendering useless an entire experimental trial. ORBIT provides a network [33] connecting the control nodes to network attached storage (NAS) and to other nodes forming a compute cluster for processing the collected data.
Figure 4: Covert communication experiment on ORBIT. Four Ettus USRP X310 radios – Alice (Tx), Bob (Rx), Willie (warden), and a broadband noise source – are linked by coaxial cables in a star topology using Mini-Circuits ZFSC-2-10G splitters/combiners. Tx-to-Rx and Tx-to-Tx path losses are 50 dB and 65 dB, respectively. All radios operate at $f_{c}=915\,\text{MHz}$ with Alice, Bob, and Willie using a DAC/ADC sampling rate of $f_{s}=12.5\times 10^{6}$ samples/s and the noise generator $f_{s}=25\times 10^{6}$ samples/s. Alice, Bob, and Willie apply 0 dB Tx/Rx gain while the noise generator applies 20 dB gain. Each X310 connects over a 10 Gb/s SFP+ link to a dedicated control node (Intel Xeon E5-2640, 20 cores); Alice’s node orchestrates the experiment via TCP messages to the other nodes and an eleven-node compute cluster of the same machines that performs real-time processing while a 2 TB network attached storage (NAS), mounted via NFS v4.2, provides shared buffer.
The radios’ internal clocks are synchronized using an Ettus OctoClock, which provides low-jitter pulse-per-second (PPS) and 10 MHz reference signals, allowing the radios to maintain a constant phase offset between each other [44]. We note that a centralized clock source is merely an experimental convenience. In practice, Alice and Bob can synchronize independently to any stable clock source, such as an atomic clock, prior to transmission. [^2] In a separate experiment in our lab at the University of Arizona, we confirmed this using a Stanford Research Systems FS725 10 MHz Rubidium Frequency Standard connected to Ettus USRP N210 radios.
Figure 5: Data-packet structure (durations in samples). A 15 400-sample preamble is created by concatenating five 13-bit Barker sequences, BPSK-modulating them, zero-padding each to 200 samples, and convolving the result with a 2 401-sample root-raised-cosine filter. The subsequent $T$-second, $T\times f_{s}$-sample, segments encode the two parts of the experimental trial. In the Alice-on segment, each shaded pulse slot begins with a strictly positive, real-valued Gaussian pilot pulse (for phase reference), immediately followed by a Gaussian-shaped, QPSK-modulated data pulse carrying randomly generated information bits. The indices of the occupied slots are drawn uniformly at random and stored in $\vec{t}$ (see Section II-C). In the subsequent Alice-off segment, she remains silent.
Alice’s node generates fresh pre-shared secret and message vectors $\vec{t}$ and $\vec{x}_{2n_{t}}$ for each experimental trial. Bob and Willies’ radios sample the channel continuously while Alice supplements the RF trial packet described in Section III-D with TCP control messages over the SFP+ connections that mark its precise start and stop times. At the end of each trial, control nodes write their data to the network-attached storage (NAS): Alice’s node stores $\vec{t}$ and $\vec{x}_{2n_{t}}$, whereas Bob and Willies’ nodes write their raw IQ samples, padded with brief pre- and post-buffers. Alice’s node then alerts the compute cluster, which analyzes the new trial, and deletes its data on completion. Acting as a rolling buffer, the NAS keeps heavy data traffic off the radio-control links and holds disk usage well below its 2 TB capacity. The Ettus USRP library and USRP hardware driver (UHD), [45, 46] are used to interact with the radios.
III-B Spectral Footprint of Alice’s Transmissions
Our analysis in Section II-D assumes that Alice’s signal energy is confined to a single $12.5$ MHz channel. In practice, this is often violated: time-limited signals are unbounded in frequency domain, and transmitter hardware non-idealities can generate strong spectral artifacts.
To characterize the emitted spectrum, we capture Alice’s waveform with a Tektronix RSA5115B real-time spectrum analyzer positioned at Willie’s location so that it records exactly what the warden observes. The analyzer is centered at $915$ MHz, samples at $2\times 10^{8}$ samples/s, and the external AWGN generator remains disabled throughout these spectral measurements. Alice transmits the shortest Gaussian pulses permitted by the system, $n_{s}^{(p)}=n_{s}^{(q)}=6$, at an amplitude of $m_{p}=0.12351232,m_{q}=1.3313578$.
Figure 6: Measured spectrum at Willie’s receiver. The Gaussian main lobe, expected to be centered at 915 MHz, is displaced by approximately +300+300 Hz, indicating a small carrier–frequency offset.
Fig. 6 plots the spectrum obtained by integrating over many successive pulses; the Gaussian main lobe is shifted by roughly +300+300 Hz from the nominal 915 MHz. This offset exceeds the datasheet specification for a GPS-disciplined 10 MHz reference, which claims a frequency accuracy of $\pm 25$ parts per billion (ppb), or $\pm 23$ Hz at 915 MHz [47].
Figure 7a overlays the pilot–data pulse pair at the transmitter with the corresponding waveform captured by the spectrum analyzer; the traces align closely, confirming that the temporal shape is maintained. Figure 7b presents the received spectrum over a wide frequency span and includes an inset which zooms to the carrier band observable at the transmitter. The transmit and receive spectra are overlaid in the inset and coincide nearly identically, confirming that the Gaussian envelope and its 99.9% energy bandwidth are maintained. The full-span view further shows that essentially all energy is confined to this zoomed region; only a few narrow spikes appear in the immediate vicinity of the band edges, and with a single capture it is unclear whether these are random artifacts or a repeatable spectral feature
Farther out in frequency, Figure 8a shows two pronounced spurs at the third and fifth harmonics of the carrier. Odd harmonics typically arise when a nominally sinusoidal signal develops a more square-like waveform, which in a transmitter can happen either through power-amplifier compression or through non-sinusoidal drive in the local oscillator (LO) chain. Neither explanation fits neatly here: the measurement is made at 0 dB hardware gain, well below the power-amplifier’s compression point, and the UBX-160 LO path is followed by a 2.2 GHz low-pass section that should strongly attenuate components at 2.75 GHz and 4.58 GHz. The precise mechanism that leaves these third- and fifth-harmonic spurs above the noise floor is therefore still under investigation.
(a) Transmitted and received pulse pair (time domain)
(b) Transmitted and received spectra
Figure 7: Carrier-band verification of the 12-sample pulse pair transmitted by Alice at 0 dB gain. Each pair comprises a 6-sample Gaussian pilot followed immediately by a 6-sample Gaussian QPSK-data pulse. The left panel overlays the time-domain waveform at the transmitter (TX) and at a spectrum-analyzer (RX) placed at Willie, while the right panel shows the RX spectra, with insets overlaying the TX and RX spectra in the carrier band.
(a) Repeated pilot–data pulse pair
(b) All-zero waveform (silent mode)gain
Figure 8: Spectra at the first five harmonics, measured at Willie with Alice’s transmitter gain set to 0 dB. Left: when Alice continuously repeats the pilot–data pulse pair. Right: when the baseband input is forced to all zeros (silent mode). The ordinate scale is re-ranged from one harmonic panel to the next to keep weak spurs visible.
To verify that Alice’s silent mode is truly quiet, we transmit an all-zero waveform. Even with an all-zero signal, the spectrum shows a small spur at the carrier and faint harmonics (Fig. 8b), behavior attributable to LO leakage.
The two spectral artifacts most relevant to covertness at the settings used in our experiments are (i) the power that spills into odd harmonics and (ii) the residual carrier spur that appears when Alice is commanded to remain silent. The former could be mitigated by inserting an external low-pass filter at Alice’s RF output, while the latter may be reduced through a careful power calibration of the X310 transmit chain. While making these improvements is challenging on a shared-testbed hardware with limited on-site access, we will pursue this in the follow-on work.
III-C Pulse Optimization
To maximize the number of covert bits that can be reliably conveyed, we jointly tune four pulse-design parameters: the pilot-pulse length $n^{(p)}_{s}$, data-pulse length $n^{(q)}_{s}$, pilot-pulse magnitude $c_{p}$, and data-pulse magnitude $c_{q}$. We model the relationship between these parameters and the attainable covert bit count as a black-box objective, apply Bayesian optimization to navigate the parameter space efficiently, and use our hardware testbed to measure the covert bits delivered for each candidate configuration.
Bayesian optimization iteratively explores the four-dimensional design space $\vec{d}=\left(n^{(p)}_{s},n^{(q)}_{s},c_{p},c_{q}\right)$ by maintaining a surrogate model of the unknown mapping from $\vec{d}$ to attainable covert bit count. We leverage Ax library in python [48] to perform this optimization. The three steps that follow are the standard Bayesian-optimization loop [49, Ch. 31]; we restate them here.
  • Propose : a new configuration $\vec{d}_{u}$ is generated from the surrogate using all observations collected so far.
  • Measure : using $\vec{d}_{u}$, the radio transmits a modified version of the data packet (Section III-D): every fifth pulse slot in the Alice-on segment carries a random bit and the Alice-off segment is removed. Number of attainable covert bits is recorded over $100$ one-second trials. The average number of attainable covert bits for a given parameter set is $r=2\,C_{\mathrm{bsc}}\,n_{p}\,\alpha_{n}$ bits over $n$ uses of the channel. In this experiment, $n$ corresponds to the number of samples available in a one second transmission. This experiment yields a sample mean $\bar{r}_{u}$ and a standard error $s_{\bar{r}_{u}}$.
  • Update : the tuple $(\vec{d}_{u},\bar{r}_{u},s_{\bar{r}_{u}})$ is appended to the data set and the surrogate is refitted before the next proposal.
After 200 iterations we report $\vec{d}^{\star}=\operatorname*{arg\,max}_{1\leq l\leq 200}\bar{r}_{l},$ the configuration that achieved the highest observed mean throughput.
We use a Gaussian-process surrogate – commonly adopted in Bayesian optimization – and briefly restate only its essentials here. For a set of $l$ different input parameter configurations $\vec{d}_{1:l}=[\vec{d}_{1},\vec{d}_{2},...\vec{d}_{l}]$, the model predicts $l$ corresponding covert throughput values sampled from the distribution $\mathcal{GP}\left(0,\,k(\vec{d}_{1:l},\vec{d}_{1:l})\right)$ [50], where $k$ is a covariance function, which in our case, is set to the Matérn-5/2 function [50, Ch. 4]. This model can also be used to predict the throughput given priors, i.e. given u1u-1 pairs of $(\vec{d},\bar{r},s_{\bar{r}})$, the model can predict the throughput $\hat{r}_{u}$ for some new parameter set $\vec{d}_{u}$. This model prediction is distributed as $\hat{r}_{u}\mid\bar{r}_{1:(u-1)},\,\vec{d}_{1:u},\,s_{\bar{r}_{1:(u-1)}}\sim\mathcal{N}(\mu_{u},\sigma_{u}^{2})$, where $\mu_{u}$ and $\sigma_{u}$ can be directly calculated. At iteration $u$ of the Bayesian optimization, configuration $d_{u}$ that maximizes expected improvement is proposed for measurement. Expected improvement for a candidate parameter configuration $d_{u}^{\prime}$ is defined as $\mathbb{E}[\hat{r}_{u}^{\prime}\mid\hat{r}_{u}^{\prime}>\bar{r}^{\star}_{1:(u-1)}]$, where $\hat{r}^{\star}_{1:(u-1)}$ is the highest measured throughput for the first u1u-1 trials, i.e. $r^{\star}_{1:(u-1)}=\max_{1\leq v\leq(u-1)}\bar{r}_{v}$
The resulting optimum $\left(n^{(p)}_{s},n^{(q)}_{s},c_{p},c_{q}\right)$ is adopted for all experiments reported henceforth.
III-D Transmission Structure for Experimental Trials
In each experimental trial, Alice transmits the three-segment packet shown in Fig. 5: a non-covert preamble, an Alice-on segment, and an Alice-off (noise-only) segment. An additive-white-Gaussian-noise (AWGN) generator operates continuously throughout the entire experiment. The Alice-on and Alice-off segments each last $T$ seconds, while the preamble marks the start of every trial and facilitates synchronization at Bob and Willie.
The preamble consists of the 13-bit Barker sequence repeated five times (65 bits in total). Each bit is BPSK-modulated and occupies 200 samples; the first sample carrying the modulated symbol followed by 199 zero-padding samples. The resulting 13 000-sample sequence is then pulse-shaped with a root-raised-cosine (RRC) filter with roll-off factor $\beta=0.35$ spanning 12 symbols, represented digitally by $12\times 200+1=2401$ taps. After filtering, the preamble length is $65\times 200+(2401-1)=15,400$ samples.
During the Alice-on segment, Alice embeds her covert message using the pilot- and data-pulse pairs defined in Section 2, transmitting only in the pulse slots indexed by $\vec{t}$. We evaluate covert throughput and Bob’s decoding reliability, as well as Willie’s detection performance under hypothesis $H_{1}$, using the Alice-on segment. The Alice-off segment, which contains only AWGN, is used to characterize Willie’s detector under the null hypothesis $H_{0}$.
III-E Experiment design
We select eight linearly-spaced values of Alice’s transmission duration $T\in[1,8]$ s. For each value of $T$, we conduct $N=150$ independent trials using distinct transmission packets described in Section III-D. Before we begin our experiment, we estimate Willie’s SNR using a $500$ 0.5 second calibration transmission with a modified version of the data-packet described in subsection III-D: every fifth pulse slot in the Alice-on segment is used to transmit a random bit and Alice-off segment is deleted. The SNR is computed from the mean power values of the pilot, pulse, and noise, and the full estimator is detailed in Appendix C. We use the SNR estimate to compute $\zeta=\alpha_{n}\sqrt{n}=\left(\mathrm{SNR}\sqrt{1+r_{p/q}^{4}}\right)^{-1}\sqrt{32\log\left(\frac{1}{1-2\delta^{2}}\right)n_{s}}$. We set $\delta=0.05$ and use the result to compute $\alpha_{n}$ for each value of $T$. This is used to generate random vectors $\vec{t}$ with pulse locations used by Alice to transmit for each trial.
III-F Results
Fig. 9 plots Bob’s receiver’s performance vs. transmission duration $T$. Bob uses $\vec{t}$ to estimate bits only in the pulse locations that Alice uses for transmission, per sparse coding described in Section II-C. We report the decoding error probability $p_{e,\mathrm{bsc}}^{(b)}$ estimated by averaging over $N=150$ trials using the right ordinate of Fig. 9. We observe that $p_{e,\mathrm{bsc}}^{(b)}\approx 0.455$ throughout our experiments. Using the left ordinate we report the corresponding estimate of the total number of transmissible covert bits $B_{\rm bsc}(n)$ using the equality in (8). Fitting a line with slope of one-half to the log-log plot results in the coefficient of determination $R^{2}=0.91$, indicating the SRL-scaling that we expect.
Fig. 10 presents Willie’s detector performance. We employ the SNR estimation strategy described in Appendix C and plot using the right ordinate. We then obtain the lower bound on Willie’s probability of error $p_{e}^{(w)}$ by computing the upper bound on Hellinger distance between the probability distributions of his received signals under each hypothesis in Appendix A. We plot this estimate using the left ordinate and note that it is very conservative compared to the covertness requirement $p_{e}^{(w)}\geq 1-\delta=0.45$ as this is, effectively, a lower bound (11) on a lower bound (9). Because our initial SNR estimate closely matches the value measured during the experiments as $T$ increases, Alice can transmit while keeping the adversary’s error probability close to the desired level of $p_{e}^{(w)}=0.45$. This empirically demonstrates the viability of the SRL-governed sparse signaling approach.
Figure 9: Bob’s receiver performance. Total number $B_{\mathrm{bsc}}(n)$ of reliably-decodable bits (left ordinate) and decoding error probability $p_{e,\mathrm{bsc}}^{(b)}$ (right ordinate) are plotted vs. transmission duration $T$, with 95% confidence intervals as error bars given $N=150$ trials per datapoint. Abscissa and left ordinate use a logarithmic scale; right ordinate uses a linear scale. Number of channel uses is $n=f_{s}T$.
Figure 10: Willie’s detector performance. Detection probability error $p_{e}^{(w)}$ (left ordinate) and SNR in dB (right ordinate) are plotted vs. transmission duration $T$, with 95% confidence intervals shown as error bars given $N=150$ trials per datapoint. Note that the left ordinate ranges from 0.40 to 0.475, and the right ranges from -20 dB to -50 dB. The abscissa is on a logarithmic scale with $\mathrm{SNR}\triangleq\frac{h_{w}^{2}c_{q}^{2}}{\sigma^{2}_{w}}$
IV Conclusion, Discussion, and Future Work
We demonstrate the first implementation of SRL-based covert communication in RF domain. We employ SDRs, which are not specifically designed for covert communication. We have to address significant challenges:
  • Dynamic range and ADC/DAC granularity limitations: The UBX-160 daughterboards use a 16-bit DAC for transmission and a 14-bit ADC for reception. While Alice can generate high-resolution low-power pulses, Bob’s lower ADC precision limits his ability to distinguish weak signals from noise. Ensuring detection of pulses after digitization leads to minimum transmission power, requiring the use of sparse coding. This, in turn, needs tight time synchronization, which we demonstrate.
  • Time and frequency synchronization: The oven-controlled crystal oscillators (OCXOs) in USRP X310s lack the frequency accuracy and stability needed to decode sparsely transmitted covert symbols. Furthermore, Bob has to know precisely when the Alice begins transmitting. We utilize a common reference (OctoClock) and a non-covert preamble, however, in practice, GPS or a highly stable time source such as atomic clock can be used for both disciplining the local oscillators and timing information. We will employ these in follow-on experiments.
  • Data volume: Our capture for $N=150$ trials amounts to approximately $3$ h of baseband data at a sampling rate of $f_{s}=12.5$ MHz, using 64-bit in-phase and quadrature samples for each receiver (Bob and Willie). This produces roughly $400$ GB of data. Although real-time processing is performed on the compute cluster, both computational and storage demands must be reduced for practical deployment.
  • Decoding performance degrades for longer durations: Transmissions with duration $T>4$ suffer from significant increase in Bob’s decoding error probability. Possible causes include LO instability and phase errors that our scheme does not correct. We will investigate means to mitigate this in future work.
  • Continuous noise injection: Maintaining an always-on, spectrally flat artificial noise floor is limited by the maximum DAC sampling rate the system can sustain: sampling too fast can overwhelm the host CPU and lead to buffer under-flows. In the future, we will explore using more than one radio to emulate noise as well as more natural noise sources by experimenting “in the wild.”
Indeed, our proof-of-concept experiment raises many theoretical and experimental research questions. In the short term, we plan to address some of them by improving our ORBIT-based design as follows:
  • We will add control over path loss to study the impact of channel dynamics;
  • We will estimate Willie’s detection error probability directly by estimating the output distributions for the optimal test statistics, as is done in [20].
  • We will investigate the optimal constellation satisfying the Hellinger distance covertness criterion. There is evidence for QPSK being such constellation, as it is optimal for complex-valued circularly-symmetric AWGN under relative entropy criterion (by extension from the BPSK case for real-valued AWGN channels examined in [2, 9, 10]). We note that this is an active topic of investigation in quantum information theory [27].
Additionally, AWGN channel model provides only a zeroth-order approximation to practical RF channels. Thus, SRL-based covert communication needs to be validated in a more realistic, dynamic environment. We plan on evolving our system to be independent rather than centrally-controlled. This would allow not only experimentation “in the wild” but also exploration of covert networks. Furthermore, we plan on relaxing assumption on adversary’s capabilities and studying the impact of, e.g., lack of precise knowledge of the timing of the transmission and pulse shape used.
Acknowledgment
The authors are grateful to Ivan Seskar and ORBIT staff for setting up the radios on ORBIT. The authors thank Loukas Lazos, Ming Li, Jingcheng Li, Ziqi Xu, Samuel H. Knarr, and Timothy C. Burt for valuable advice, and sharing the equipment. Finally, the authors acknowledge helpful discussions with Mark J. Meisner, Jaim Bucay, Dennis L. Goeckel, Donald F. Towsley, Matthieu R. Bloch, Matthew Arcarese, and Robert J. McGurrin.
Appendix A Covertness Criterion Analysis
[…]
$\psi(\vec{w};\vec{\mu},\mathbf{\Gamma},\mathbf{C})={\frac{\exp\!\left(-{\frac{1}{2}}{\begin{bmatrix}\vec{w}-\vec{\mu}\\ {\overline{\vec{w}}}-{\overline{\vec{\mu}}}\end{bmatrix}}^{\dagger}{\begin{bmatrix}\mathbf{\Gamma}&\mathbf{C}\\ {\overline{\mathbf{C}}}&{\overline{\mathbf{\Gamma}}}\end{bmatrix}}^{\!\!-1}\!{\begin{bmatrix}\vec{w}-\vec{\mu}\\ {\overline{\vec{w}}}-{\overline{\vec{\mu}}}\end{bmatrix}}\right)}{\pi^{n}{\sqrt{\det(\mathbf{\Gamma})\det(\overline{\mathbf{\Gamma}}-\mathbf{C}^{\dagger}\mathbf{\Gamma}^{-1}\mathbf{C})}}}}.$ (14)
Since our noise is i.i.d. and induces circularly symmetric complex Gaussian distributions, $\mathbf{C}=\mathbf{0}$ and we drop it for brevity. Under $H_{0}$, Alice does not transmit and Willie observes AWGN. Thus, segments are i.i.d. with the density function:
$p^{(0)}\left(\vec{w}_{i}\right)$ $=\psi\left(\vec{w}_{i};\vec{0},2\sigma^{2}_{w}\mathbf{I}_{n_{s}}\right).$ (15)
Alice transmits under hypothesis $H_{1}$. Since Willie does not have $\vec{t}$ and $\vec{s}$, segments are i.i.d. with the density function:
$p^{(1)}\left(\vec{w}_{i}\right)$ $=\int_{0}^{2\pi}\mathop{}\!\mathrm{d}\theta\left((1-\alpha_{n})\psi\left(\vec{w}_{i};\vec{0},2\sigma^{2}_{w}\mathbf{I}_{n_{s}}\right)\right$.
$\phantom{=\int==}\left.+\frac{\alpha_{n}}{4}\sum_{x=0}^{3}\psi\left(\vec{w}_{i};e^{j\theta}\vec{\mu}(x),2\sigma^{2}_{w}\mathbf{I}_{n_{s}}\right)\right)$
$=(1-\alpha_{n})\psi\left(\vec{w}_{i};\vec{0},2\sigma^{2}_{w}\mathbf{I}_{n_{s}}\right)$ (16)
$\phantom{=}+\frac{\alpha_{n}}{4}\sum_{x=0}^{3}\int_{0}^{2\pi}\mathop{}\!\mathrm{d}\theta\psi\left(\vec{p}_{i};e^{j\theta}h_{w}\vec{c}_{p},2\sigma^{2}_{w}\mathbf{I}_{n_{s}}\right)$
[…]
$=(1-\alpha_{n})\psi\left(\vec{w}_{i};\vec{0},2\sigma^{2}_{w}\mathbf{I}_{n_{s}}\right)$
$\phantom{=}+\frac{\alpha_{n}}{4}\psi\left(\vec{w}_{i};\vec{0},2\sigma^{2}_{w}\mathbf{I}_{n_{s}}\right)e^{-a_{w}^{2}\frac{\|\vec{c}_{p}\|^{2}+\|\vec{c}_{q}\|^{2}}{2\sigma_{w}^{2}}}$ (18)
$\phantom{=+}\times\sum_{x=0}^{3}I_{0}\left(\frac{a_{w}}{\sigma_{w}^{2}}f_{x}\left(\vec{w}_{i},\vec{c}_{p},\vec{c}_{q}\right)\right),$ (19)
where $I_{0}(\cdot)$ is the modified Bessel function of the first kind with order 0, $a_{w}$ is the channel loss, and
$f_{k}$ $\left(\vec{w}_{i},\vec{c}_{p},\vec{c}_{q}\right)$
$=\left(\left(\left\langle\vec{c}_{p},\Re(\vec{p}_{i})\right\rangle+(-1)^{\lfloor k/2\rfloor}\left\langle\vec{c}_{q},\Re(\vec{q}_{i})\right\rangle\right)^{2}\right$.
[…]
We seek to bound Willie’s probability of detection error. Cauchy-Schwarz inequality implies $\mathcal{V}_{T}(P_{0}^{n},P_{1}^{n})\leq\sqrt{2}H(P_{0}^{n},P_{1}^{n})$, thus we analyze the Hellinger distance between our distributions. We have
$H^{2}$ $(P_{0}^{n},P_{1}^{n})$
$=1-\int_{\mathbb{C}^{n}}\mathop{}\!\mathrm{d}^{2}\vec{w}_{i}\sqrt{\prod_{i=1}^{n_{p}}p^{(0)}\left(\vec{w}_{i}\right)p^{(1)}\left(\vec{w}_{i}\right)}$ (21)
$=1-\prod_{i=1}^{n_{p}}\int_{\mathbb{C}^{n_{s}}}\mathop{}\!\mathrm{d}^{2}\vec{w}_{i}\sqrt{p^{(0)}\left(\vec{w}_{i}\right)p^{(1)}\left(\vec{w}_{i}\right)}$ (22)
$=1-\left(\int_{\mathbb{C}^{n_{s}}}\mathop{}\!\mathrm{d}^{2}\vec{w}_{1}\sqrt{p^{(0)}\left(\vec{w}_{1}\right)p^{(1)}\left(\vec{w}_{1}\right)}\right)^{n_{p}}$ (23)
The term inside the power of $n_{p}$ on the right hand side of (23) admits no known closed-form expression, so we take its Taylor series expansion. Without loss of generality, substitute $\vec{c}_{q}\to c_{q}\vec{b}_{q}$ where $c_{q}=\|\vec{c}_{q}\|$ is the magnitude of the QPSK pulse and $\|\vec{b}_{q}\|=1$ is the unit vector with the same pulse shape as $\vec{c}_{q}$. Similarly, substitute $\vec{c}_{p}\to c_{q}r_{p/q}\vec{b}_{p}$ where $r_{p/q}=\frac{\|\vec{c}_{p}\|}{\|\vec{c}_{q}\|}$ is the ratio of magnitudes the pilot pulse and the QPSK pulse, and $\|\vec{b}_{p}\|=1$ is the unit vector with the same pulse shape as the $\vec{c}_{p}$. We Taylor expand the term inside the power of $n_{p}$ about $c_{q}=0$, and obtain:
$H^{2}$ $(P_{0}^{n},P_{1}^{n})$
$=1-\alpha_{n}^{2}\frac{1+r_{p/q}^{4}}{32\sigma_{w}^{4}}a_{w}^{4}c_{q}^{4}+\alpha_{n}^{3}\frac{1+r_{p/q}^{6}}{64\sigma_{w}^{6}}a_{w}^{6}c_{q}^{6}+o(c_{q}^{6})$ (24)
Taylor’s theorem with remainder implies
$H^{2}(P_{0}^{n},P_{1}^{n})$ $\leq 1-\left(1-\alpha_{n}^{2}\frac{1+r_{p/q}^{4}}{32\sigma_{w}^{4}}a_{w}^{4}c_{q}^{4}\right)^{n_{p}}.$ (25)
We seek to bound $\alpha_{n}$ to ensure covertness. Using covertness requirement $\delta$ and (11) implies
$\delta\leq\sqrt{\frac{1}{2}\left(1-\left(1-\alpha_{n}^{2}\frac{1+r_{p/q}^{4}}{32\sigma_{w}^{4}}a_{w}^{4}c_{q}^{4}\right)^{n_{p}}\right)}$ (26)
Rearranging yields
$\alpha_{n}\leq\frac{4\sigma_{w}^{2}}{a_{w}^{2}c_{q}^{2}}\sqrt{\frac{2\left(1-(1-2\delta^{2})^{1/n_{p}}\right)}{1+r_{p/q}^{4}}}.$ (27)
Now, we substitute $1/n_{p}\to x^{2}$ and Taylor expand the right-hand side of (27) with respect to $x=0$ as
$\frac{4\sigma_{w}^{2}}{a_{w}^{2}c_{q}^{2}}\sqrt{\frac{2\left(1-(1-2\delta^{2})^{x^{2}}\right)}{1+r_{p/q}^{4}}}$
$\phantom{==}=\frac{4\sigma_{w}^{2}}{a_{w}^{2}c_{q}^{2}}\sqrt{\frac{2\log\left(\frac{1}{1-2\delta^{2}}\right)}{1+r_{p/q}^{4}}}x$
$\phantom{===}-\frac{\sigma_{w}^{2}}{a_{w}^{2}c_{q}^{2}}\sqrt{\frac{2\log\left(\frac{1}{1-2\delta^{2}}\right)}{1+r_{p/q}^{4}}}\log\left(\frac{1}{1-2\delta^{2}}\right)x^{3}+o(x^{3}).$ (28)
Then, for $n_{p}$ large enough, Taylor’s theorem with remainder and reverting the substitution implies
$\alpha_{n}\leq\frac{4\sigma_{w}^{2}}{a_{w}^{2}c_{q}^{2}}\sqrt{\frac{2\log\left(\frac{1}{1-2\delta^{2}}\right)}{n_{p}(1+r_{p/q}^{4})}}.$ (29)
Appendix B Phase Estimation
In each pilot segment of the pulse, Bob receives the complex-valued vector
$\vec{y}_{p}$ $=a_{b}\vec{c}_{p}e^{j\theta_{b}}+\vec{z}^{(b)}_{p}$, (30)
where $\vec{z}^{\,(b)}_{p}\in\mathbb{C}^{n_{s}^{(p)}}$ is circularly-symmetric complex-valued AWGN afflicting the pilot segment. Applying the pilot-pulse-shape filter yields:
$\langle\vec{c}_{p},\vec{y}_{p}\rangle$ $=a_{b}e^{j\theta_{b}}\|\vec{c}_{p}\|^{2}+\langle\vec{c}_{p},\vec{z}^{(b)}_{p}\rangle$. (31)
The expected values of the in-phase and quadrature (IQ) components $p_{I}\triangleq\Re\left(\langle\vec{c}_{p},\vec{y}_{p}\rangle\right)$ and $p_{Q}\triangleq\Im\left(\langle\vec{c}_{p},\vec{y}_{p}\rangle\right)$ are:
$\mathbb{E}\left[\Re\left(\langle\vec{c}_{p},\vec{y}_{p}\rangle\right)\right]$ $=\mathbb{E}[p_{I}]=a_{b}\|\vec{c}_{p}\|^{2}\cos(\theta_{b})$ (32)
$\mathbb{E}\left[\Im\left(\langle\vec{c}_{p},\vec{y}_{p}\rangle\right)\right]$ $=\mathbb{E}[p_{Q}]=a_{b}\|\vec{c}_{p}\|^{2}\sin(\theta_{b}),$ (33)
since AWGN is circularly-symmetric and has zero mean. The estimate of phase $\theta_{b}$ is then $\hat{\theta}_{b}=\tan^{-1}\frac{p_{I}}{p_{Q}}$. As $n_{s}^{(p)}$ increases, $p_{I}$ and $p_{Q}$ converge to their expected values, yielding a more accurate estimate of phase. Here, $n_{s}^{(p)}=6$ suffices.
Appendix C Estimation of Willie’s SNR
We need to estimate Willie’s SNR given the knowledge of $\vec{c}_{p}$, $\vec{c}_{q}$, as well as transmitted symbols in $\vec{x}_{n_{t}}$ and their locations $\vec{t}$. While Willie has no access to $\vec{x}_{n_{t}}$ and $\vec{t}$, we use them to characterize the SNR of his system. A pulse-bearing slot received by Willie is:
$\vec{w}_{p}(x)$ $=a_{w}e^{j\theta_{w}}\vec{u}(x)+\vec{z}^{(w)},$ (34)
for $x\in\mathcal{X}$ and unknown phase $\theta_{w}$. Applying the pulse shape $\vec{u}(x)$ corresponding to transmitted $x$ as a filter yields:
$\langle\vec{u}(x),\vec{w}_{p}(x)\rangle$ $=a_{w}e^{j\theta_{w}}c^{2}+\langle\vec{u}(x),\vec{z}^{(w)}\rangle,$ (35)
since $\|\vec{u}(x)\|=c$ per Section 2. The expected value of the above is:
$\mathbb{E}\left[\langle\vec{u}(x),\vec{w}_{p}(x)\rangle\right]$ $=a_{w}e^{j\theta_{w}}c^{2},$ (36)
as AWGN has zero mean, and, hence, $\mathbb{E}\left[\langle\vec{u}(x),\vec{z}^{(w)}\rangle\right]=0$. Thus, averaging over many pulse-bearing slot observations, and dividing by a known constant $c^{2}$, yields an estimate of $a_{w}e^{j\theta_{w}}$. Its squared magnitude is the estimate $\hat{a}_{w}^{2}$ of $a_{w}^{2}$.
An empty pulse slot received by Willie contains only noise:
$\vec{w}_{p}(0)$ $=\vec{z}^{(w)},$ (37)
The expectation of its squared magnitude is:
$\mathbb{E}\left[\langle\vec{w}_{p}(0),\vec{w}_{p}(0)\rangle\right]$ $=\mathbb{E}\left[\langle\vec{z}^{(w)},\vec{z}^{(w)}\rangle\right]=2n_{s}\sigma_{w}^{2}.$ (38)
Thus, averaging over many instances of empty pulse slots, and dividing by a known constant $2n_{s}$, yields an estimate $\hat{\sigma}_{w}^{2}$ of $\sigma_{w}^{2}$. Finally, we estimate the SNR by evaluating $\frac{\hat{a}_{w}^{2}c_{q}^{2}}{\hat{\sigma}_{w}^{2}}$, for known $c_{q}=\|\vec{c}_{q}\|$.